1. Differentiate of the following functions with respect to x:

i) 𝑆𝑖𝑛π‘₯βˆ™ ln(5π‘₯)

ii) βˆšπ‘π‘œπ‘ βˆšπ‘₯

iii) 𝑒 ln (π‘‘π‘Žπ‘›5π‘₯)

iv) 𝑆𝑖𝑛2 {ln(𝑠𝑒𝑐π‘₯)}

v) ln (π‘‘π‘Žπ‘›π‘₯) οΏ½


1
Expert's answer
2021-10-26T10:39:01-0400

"(\sin(x))'=\cos(x)"

"(\cos(x))'=-\sin(x)"

"(\ln(x))'=\frac{1}{x}"

"(\sqrt{x})'=\frac{1}{2}\cdot\frac{1}{\sqrt{x}}"

"(e^{x})'=e^{x}"

"(\tan(x))'=\frac{1}{\cos^2(x)}"

"(C\cdot x)'=C"


Let, "h(x)=f(g(x)), h'(x)=f'(x)\cdot g'(x)" (*)

Use(*),

"(\sec(x))'=(\frac{1}{\cos(x)})'=(-\frac{1}{\cos^{2}(x)})\cdot(-\sin(x))=\frac{\sin(x)}{\cos^2(x)}" ,

where "f(x)=\frac{1}{\cos(x)}, g(x)=\cos(x)" ).

-------------------------------------------------------------------------------------------------------------------

1) "(\sin(x)\cdot\ln(5x))'."

"(f\cdot g)'=f'g+fg'" (1)

Use (1),

"(\sin(x)\cdot\ln(5x))'=(\sin(x))'\cdot\ln(5x)+\sin(x)\cdot(\ln(5x))'="

"=\cos(x)\cdot\ln(5x)+\sin(x)\cdot\frac{1}{5x}\cdot5=\cos(x)\cdot\ln(5x)+\frac{\sin(x)}{x}".

-------------------------------------------------------------------------------------------------------------------

2) "(\sqrt{\cos(\sqrt{x})})'".

Let "p(x)=f(g(h(x)))", then "p'(x)=f'(x)\cdot g'(x)\cdot h'(x)".

"f(x)=\sqrt{u}, u=\cos(\sqrt{x})\implies f'(x)=\frac{1}{2}\cdot\frac{1}{\sqrt{u}}=\frac{1}{2}\cdot\frac{1}{\sqrt{\cos(\sqrt{x})}}",

"g(x)=\cos(v),v=\sqrt{x}\implies g'(x)=-\sin(v)=-\sin(\sqrt{x})",

"h(x)=\sqrt{x}, h'(x)=\frac{1}{2}\cdot\frac{1}{\sqrt{x}}".

"(\sqrt{\cos(\sqrt{x})})'=\frac{1}{2}\cdot\frac{1}{\sqrt{\cos(\sqrt{x})}}\cdot(-\sin(\sqrt{x}))\cdot\frac{1}{2}\cdot\frac{1}{\sqrt{x}}="

"=-\frac{1}{4}\cdot\frac{\sin(\sqrt{x})}{\sqrt{x}\cdot\sqrt{\cos(\sqrt{x})}}".

-------------------------------------------------------------------------------------------------------------------

3) "(e^{\ln(\tan(5x))})'".

Let "p(x)=s(f(g(h(x))))\implies p'(x)=s'(x)\cdot f'(x)\cdot g'(x)\cdot h'(x)".

"s(x)=e^{u}, u=\ln(\tan(5\cdot x))\implies s'(x)=e^{u}=e^{\ln(\tan(5\cdot x))}",

"f(x)=\ln(v), v=\tan(5\cdot x)\implies f'(x)=\frac{1}{v}=\frac{1}{\tan(5\cdot x)}",

"g(x)=\tan(w), w=5\cdot x\implies g'(x)=\frac{1}{\cos^{2}(w)}=\frac{1}{\cos^2(5\cdot x)}",

"h(x)=5\cdot x\implies h'(x)=5".

"(e^{\ln(\tan(5x))})'=e^{\ln(\tan(5x))}\cdot\frac{1}{tan(5x)}\cdot\frac{1}{\cos^{2}(5x)}\cdot5="

"=5\cdot\tan(5x)\cdot\frac{1}{\tan(5x)}\cdot\frac{1}{\cos^{2}(5x)}=5\sec^{2}(x)".

-------------------------------------------------------------------------------------------------------------------

4) "(\sin(2(\ln(\sec(x)))))'".

"(\sin(2(\ln(\sec(x)))))'=\cos(2(\ln(\sec(x)))))\cdot2\cdot\frac{1}{\sec(x)}\cdot\frac{\sin(x)}{\cos^2(x)}="

"=2\cdot\cos(2(\ln(\sec(x)))))\cdot\tan(x)".

-------------------------------------------------------------------------------------------------------------------

5) "(\ln(\tan(x)))'=\frac{1}{\tan(x)}\cdot\frac{1}{\cos^2(x)}=\frac{1}{\sin(x)}=\csc(x)".